OpenStudy (anonymous):

how do you factor the trinomial, please explain (ex: x^2 +8x+7

OpenStudy (anonymous):

( x+7) (x+1)

OpenStudy (anonymous):

therefore x+7 =0 , where X= -7 and x+1 = 0, where x = -1

OpenStudy (amistre64):

the "how" might help you factor others tho ..... since of course they dont ALL factor into (x+7) and (x+1)

OpenStudy (amistre64):

if a trinomial is going to factor out, it will take the form: (px+q)(mx+n) ; expanding this out we get pm x^2 + pnx + qm x + qn pm x^2 + (pn +qm) x + qn we can use this to compare your example with by equating coefficients pm x^2 + (pn +qm) x + qn = x^2 +8x+7 pm = 1 pn+qn = 8 qn = 7

OpenStudy (amistre64):

when the example has 1 as a leading coeff, we can let p=m=1, which simplifies this to n+q = 8 qn = 7 so we are looking for factors of 7, that add up to 8 7(1) = 7, 7+1 = 8 those fit